OpenStudy (anonymous):

What is the GCF of 16 and 72? A. 16 B. 2 C. 8 D. 4

OpenStudy (whpalmer4):

Well, as you get the luxury of choosing from 1 of 4 possible answers, you could just try them, starting from the largest, and selecting the largest one that divides evenly (read: is a factor of) into both 16 and 72.

OpenStudy (whpalmer4):

To actually compute the answer: Factor both numbers into their prime factorizations: 16=2*2*2*2 = 2^4 72=2*2*2*3*3 = 2^3*3^2 Now go through and collect the highest common power of all of the common factors, and multiply them together. That's your GCF. An example: find the GCF of 54 and 72 54 = 2*3*3*3 = 2^1*3^3 72 = 2*2*2*3*3 = 2^3*3^2 Highest common power of 2 is 2^1. Highest common power of 3 is 3^2. GCF is 2^1*3^2 = 2*9 = 18.